Reports 1-1 of 1 Clear search Modify search
Virgo Runs (O4c)
berni - 23:00 Wednesday 01 October 2025 (67837) Print this report
Operator Report - Afternoon shift

ITF found in lock acquisition.

Diego investigated the locking issueswhich at the end seemed to be due mostly to the bad weather; Science mode was finally set at 17:32 UTC.

 

Guard tours (time in UTC)

18:10-18:40

Images attached to this report
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×